OpenStudy (anonymous):

Please help i have no idea how to do this: Find the inverse of the function. f(x) = 3sqrt(x/9)- 4

OpenStudy (johnweldon1993):

You switch your 'x' and 'y' and solve for 'y' again

OpenStudy (anonymous):

To solve for inverse, change the x into y and y into x and then solve for y!

OpenStudy (johnweldon1993):

\[y = 3\sqrt{\frac{ x }{ 9 }} - 4\] switch the 'x' and 'y' \[x = 3\sqrt{\frac{ y }{ 9 }} - 4\] how would you solve for 'y' again?

OpenStudy (anonymous):

i dont know how to do that with the cubed root though :/

OpenStudy (johnweldon1993):

Oh that's a cubed root? oh sorry...let me do that again lol

OpenStudy (johnweldon1993):

\[x = \sqrt[3]{\frac{ y }{ 9 }} - 4\]

OpenStudy (anonymous):

well i know you add four to both sides right? but im not sure how to continue

OpenStudy (johnweldon1993):

Well....you want to get 'y' by itself...so first add 4 to both sides.....what do you have on the left side now?

OpenStudy (johnweldon1993):

nvm lol you beat me to that part haha

OpenStudy (anonymous):

:)

OpenStudy (johnweldon1993):

Well you want to get rid of that radical....to get rid of a SQUARE root....you raise both sides to the 2nd power.... so to get rid of a CUBED root....you raise both sides to the 3rd power right?

OpenStudy (anonymous):

oh ok! so youd have 64=y/9, then y=64/9 or 7.11?

OpenStudy (anonymous):

just kidding

OpenStudy (johnweldon1993):

You forgot your 'x' On the left side ....lol nvm continue :D

OpenStudy (anonymous):

hah sorry i forgot about the x yeah. ok.. so 64x= y/9? then y=64x/9?

OpenStudy (johnweldon1993):

Not quite on the left side before you cubed both sides you have x + 4 right? so when you cube both sides you get (x + 4)^3 = y/9 then you would multiply both sides by 9 and get.....?

OpenStudy (anonymous):

oh yeah ok hang on a sec

OpenStudy (anonymous):

so (x^3+12x^2+48x+64)/9=y

OpenStudy (anonymous):

oh i think the answer is just 9(x+4)^3, sorry i kept messing up. i think i got it now.

OpenStudy (johnweldon1993):

lol there you go :) don't simplify unless you really have to :)

OpenStudy (anonymous):

thanks youre the best!

OpenStudy (johnweldon1993):

anytime :)